A ship has an area of cross section at the water line 200cm^2

a.by what depth does the ship sink in fresh water when it loads a cargo of 4000 tonnes

b. if the ship plus cargo has a displacement turnage of 12300 tonnes, by what amount will the ship raise in the water when it sails from fresh water into sea water


1
Expert's answer
2021-06-15T10:01:10-0400

a. "F= PA= \rho gh A \implies h = \frac{F}{\rho g A}=\frac{4000*10000}{200*1000*9.81} = 20.39 cm"

b. "V= \frac{m}{\rho}=\frac{12300*1000}{1029-1000}=424137.93 m^3"